OpenStudy (cookiimonster627):

Write a function with the given characteristics: a polynomial with rational coefficients having roots 3, 3, and 3-i

OpenStudy (cookiimonster627):

@Vocaloid

jimthompson5910 (jim_thompson5910):

3-i is one root the conjugate 3+i is another root they come in pairs

jimthompson5910 (jim_thompson5910):

x = 3-i or x = 3+i x-3 = -i or x-3 = i (x-3)^2 = (-i)^2 or (x-3)^2 = (i)^2 (x-3)^2 = -1

jimthompson5910 (jim_thompson5910):

if you were to solve (x-3)^2 = -1 for x, you'd get x = 3-i or x = 3+i

jimthompson5910 (jim_thompson5910):

(x-3)^2 = -1 (x-3)^2 + 1 = 0 x^2 - 6x + 9 + 1 = 0 x^2 - 6x + 10 = 0 so if 3-i is a root, then x^2 - 6x + 10 is a factor

jimthompson5910 (jim_thompson5910):

if 3 is a root, then x-3 is a factor this is a double root, so (x-3)^2 is a factor

jimthompson5910 (jim_thompson5910):

x^2 - 6x + 10 is a factor (x-3)^2 is a factor so you'll have (x-3)^2*(x^2-6x+10) expand that all out to get the final answer

OpenStudy (cookiimonster627):

what do you mean by expand that out?

OpenStudy (cookiimonster627):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hopefully you see how (x-3)^2 turns into x^2 - 6x + 9 ?

OpenStudy (cookiimonster627):

you get that by multiplying (x-3)(x-3)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

(x-3)^2*(x^2-6x+10) turns into (x^2-6x+9)*(x^2-6x+10)

jimthompson5910 (jim_thompson5910):

now expand out (x^2-6x+9)*(x^2-6x+10) you can use the box method (which is what I would do)